OpenStudy (faman39):

For a summer fundraiser, Rob is doing manual labor to help several families do home improvement projects. He is tracking the amount of time he spends each week with a spreadsheet. If his spreadsheet shows that during the past two weeks, he spent a total of 11,040 minutes working, how many days and hours has he worked? 6 days and 6 hours 6 days and 16 hours 7 days and 7 hours 7 days and 16 hours

OpenStudy (faman39):

@LukeBlueFive can you help me with this?

OpenStudy (lukebluefive):

Sure. There are 60 minutes in an hour: \[1 hour = 60 minutes\] So this means: \[11,040 minutes \times \frac{1hour}{60 minutes}\] Will give you the equivalent number of hours. Furthermore, there are 24 hours in a day: \[1 day = 24 hours\] So this means you need to multiply the previous result by: \[\frac{1 day}{24 hours}\] To get the equivalent number of days.

OpenStudy (faman39):

66240 and 264090, thats what i am getting :( so sorry

OpenStudy (lukebluefive):

You seem to be multiplying instead of dividing... guess I should've made my equations clearer. Let me rewrite them: \[\frac{11,040 minutes}{60 minutes}\] Will give you the number of hours. Then, divide the following from that result: \[24 hours\] To get the number of days.

OpenStudy (lukebluefive):

You divide since there are more minutes in a day than hours, and more hours than days in a year... if that makes any sense.

OpenStudy (faman39):

7.666667 i got final answer this

OpenStudy (faman39):

184 dividing 11040/60 184 / 24 got that 7.6666667

OpenStudy (lukebluefive):

Okay, he worked for over 7 days, but how much over? To find out, multiply the .6666667 part by 24 to get the number of hours that is equal to 2/3 of a day.

OpenStudy (lukebluefive):

Then you'll have the answer: 7 days and ? hours

OpenStudy (faman39):

But i m getting 184 when i multiply

OpenStudy (faman39):

Sorry got 16, you meant just multiply .666667 with 24

OpenStudy (faman39):

Thank you so sooooooooooo much @LukeBlueFive !!!!!!!!!! your totally amazing !!!!!

OpenStudy (faman39):

So sorry for troubling alot for this probs, actually worst at math

OpenStudy (lukebluefive):

You're too kind, but you're welcome. Hope this made sense.

OpenStudy (faman39):

yes :) thank you so much, n um so sorry again for trouble

OpenStudy (lukebluefive):

That's what I'm here for. =D

OpenStudy (faman39):

Lol thanks alot, your really great and truly genius !!! :)

OpenStudy (lukebluefive):

lol, I've just seen a lot of math over the years... ^_^;